Please answer A,b,c and D
tem 2 Part A Find the pH of each of the following solutions of mixtures of acids. 0.105 M in HBr and 0.150 M in HCHO2 Express your answer to three decilmal places. pH = Submit My Answers Give Up Part B 0.130 M in HNO2 and 9.0 102 M in HNO Express your answer to two decimal places. Submit My Answers Give Up Part C 0.190 M in HCHO2 and 0.23 M in HC2H0 Express your answer to two decimal places
media%2Fbdb%2Fbdb2159d-a2d3-48ef-bac3-a7
Answer #1

part A)

concentration of H+ in HBr = 0.105 M

concentration of H+ in HCHO2 = sqrt (Ka x C)

                                               = sqrt (1.8 x 10^-4 x 0.150)

                                               = 5.196 x 10^-3

total [H+] = 0.110 M

pH = -log [H+] = -log (0.110)

pH = 0.958

part B)

in HNO2 , [H+] = sqrt (5.6 x 10^-4 x 0.130) = 8.53 x10^-3 M

[H+] , in HNO3 = 9 x 10^-2 M

[H+] = 0.0985 M

pH = 1.006

part C :

[H+] = sqrt (Ka1 C1 + Ka2 C2)

       = sqrt (1.8 x 10^-4 x 0.190 + 1.8 x 10^-5 x 0.23)

[H+] = 6.19 x 10^-3

pH = 2.208

part D)

pH = 3.046